OpenStudy (valirium):

Solve the system by substitution. y = x^2 - 2x + 1 y = 2x + 1

OpenStudy (will.h):

Alright so we will substitute the value of the 2nd equation in the 1st equation and then we will be able to solve for x 2x + 1 = x^2 -2x + 1 Solve for x you can use famous quadratic formula but 1st set the entire terms on one side

OpenStudy (valirium):

What do you mean set the entire terms on one side? @Will.H

OpenStudy (will.h):

2x + 1 = x^2 -2x +1 Set the left side equal to 0 How? By getting rid of all terms Subtract 1 from both sides and 2x from both sides to get 0 = x^2 -4x Make sense?

OpenStudy (valirium):

Yes @Will.H

OpenStudy (will.h):

Alright. Use famous quadratic formula and find the x values There will be 2 solutions

OpenStudy (will.h):

Find them and tag me to see if you're correct

OpenStudy (valirium):

Umm. Would it be (2,5) and (1,3)? Or am I wrong? @Will.H

OpenStudy (will.h):

You're wrong.

OpenStudy (will.h):

I'll explain

OpenStudy (will.h):

X^2 -4x = 0 We can use factoring method GCF is x so X(x -4) = 0 So either x = 0 Or x-4 = 0 which equal x =4 Make sense?

OpenStudy (valirium):

It does @Will.H

OpenStudy (will.h):

Terrific. Now we will find the y values shall we?

OpenStudy (will.h):

Y = 2x +1 We found x = 0 Or x= 4 So let's find y when x = 0 first Y = 0 + 1 which equal y = 1 Now when x = 4 Y = 2(4) + 1 which equals y = 9 Make sense? That's all

OpenStudy (will.h):

So our solutions are (0, 1) and (4,9) You can ask if something is confusing you. I can explain

OpenStudy (valirium):

Oh it makes sense! Thank you! :) @Will.H

OpenStudy (will.h):

You welcome :)